OpenStudy (anonymous):

3y+2>5y+8 solve for y and there is a _ under the >

OpenStudy (anonymous):

I'll do an example for you to follow along with.\[7x+4 \ge 2x-11\]\[7x-2x+4 \ge 2x-2x-11\]\[5x+4-4 \ge -11-4\]\[5x \ge -15\]\[x \ge -3\]

OpenStudy (anonymous):

First, subtract 3y from both sides to get \[2 \ge 2y+8\]

OpenStudy (anonymous):

Then subtract 8 from both sides to get \[-6 \ge 2y\] Then divide by 2 and reverse the inequality to get \[y \le -3\]